OpenStudy (anonymous):

rewrite as a product of linear functions, and find the zeroes of polynomialx^4 + 5x^2-36

OpenStudy (cwrw238):

=(x^2 + 9)(x^2 - 4) = (x^2 + 9)(x + 2)(x - 2) x^2 + 9 is prime

OpenStudy (anonymous):

so we factor first?

OpenStudy (cwrw238):

yes there are 2 real roots x = 2 and x = -2 the other 2 roots are complex

OpenStudy (anonymous):

so they are all square roots?

OpenStudy (cwrw238):

x^2 + 9 = 0 x^2 = -9 x = 3i and -3i

OpenStudy (anonymous):

thank you.

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

where did you pull the 2 from?